I am asking this question to ensure I have proper understanding of the formulas. Please show work symbolically.

After colliding, two hockey players are moving together in the same direction at a velocity of 3 m/s. Hockey player 1 has a mass of 80 kg and hockey player 2 has a mass of 75 kg.

Scenario 1: The 80 kg hockey player pushes the 75 kg player away at 5 m/s in opposite direction of their motion. What is the velocity of the 80 kg skater after he pushes the other player?

Scenario 2: The 80 kg hockey player pushes the 75 kg player away at 5 m/s in the same direction as their motion. What is the velocity of the 80 kg skater after he pushes the other player?
Answer #1

Solution : I am considering the velocities which you have proved are in ground frame.

I will treat initial velocity with u and final with v.

Now for the first scenario.

Initial momentum = final momentum

M1u1 +m2u2 = m1v1 + m2v2

(80 +75)×3 = 75 x -5 + 80 x v2 ( v1 is negative because direction of motion is opposite to initial direction. )

So v2 = +10.5 m/sec

Now case b

80 x 3 + 75 x 3 = 75 x 5 + 80 x v2

So v2 = 1.125 m/sec